27 March 2011 UT: LUCI Science Prep

Telescope Operator: Geno Bechetti
Instrument Support: Olga Kuhn, Tim Shih
Telescope Support: Jeff Urban
SW support:

Summary

Preparations

Details

Observing Plan (copy and paste this to the nightly logs)

LUCI1 functional checkout:
  • Check system time vs our GPS, at cmd line execute: "ntpdate -q 10.144.0.211" DONE offset -0.000001 sec
  • Run /home/engineer/bin/showSupervisorStatus.sh, check if OK DONE
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered on DONE
  • Start LUCIFER Control SW and check that all is running as needed DONE
    • Restart TelescopeService if TCS has been restarted DONE
  • Open UI and check that telescope data are being updated DONE
  • Check darks (see: LUCIFER$/readnoise script)
    • O2DCR 2.0s (usually around 11.6 e-): luci.20110327.0002 and 0003 -> 11.77 e-
    • MER10 10.0s (usually around 5.5 e-): luci.20110327.0004 and 0005 -> 7.19 e-, there are bright corners at top left and bottom right
    • no mask was in use: put closed/blind in and repeat MER10 10.0s test: 6 and 7 -> 5.42 e-
    • repeat O2DCR too: 8 and 9 -> 11.5 e-*
  • Take sieve mask images in N3.75 and N1.8 to check focus gradients: *optic sieve going in: 10 is with N1.8 - definitely focus gradient. (forgot to click initialize...)
    • N3.75 in, initialize: 11 is with N3.75
    • back to N1.8, but error!! Error Setting the Camera position Motion of Motor card number 3, motor number: 2, type: "camera selection wheel"..., didn't move - initialized again and red background cleared. Now switched back to N1.8.
    • N1.8 is luci.20110327.0012.fits - still showing large focus gradient towards bottom.
  • Move calibration unit in moved in from UI
    • Check that lamps work *with LS600, N1.8, 200H+K 1.98mic:
      • Ne 2sec and 20sec do not show anything - blind filter was in (these were 13 and 14).
      • ordersep filter: 20 sec shows lines and continuum - there is light in dome. - vent doors open (15 NE on, 16 NE off)
      • with vent doors open hard to test this, but the GUI gets the signal that the lamps are on - turn when I click it.
      • Closing vent doors -
      • Ar ON 17, OFF 18
      • move central wavelength to 1.93mic AR ON 19 and OFF 20.
      • HALO1: ON 21 (20 too long - 50k counts). 5 sec, ON 22 and OFF 23.
    • Check gratings/tilts
      • 210zJHK at 2.22mic: 20sec Xe ON 24 with ordersep, change to K and 20 sec ON 25 K, OFF 26
    • Take an imaging flat and check levels
      • HALO2: ON 27 and OFF 28 and finally HALO3: ON 29 and OFF 30
    • Move calibration unit out moved out from UI
  • Check mask tables after exchange Tim did this - restart IMGUI and table looks OK (not yet checked by moving in/out masks)
  • Run LUCI1 flexure tests, see: ~/LBTO/Flexures/RotatorScript.txt

Other Tests: Closed dome SW tests, if requested

Twilight flats in NB, BB filters (cycle through all over time)
  • FeII w N1.8 #31 completely saturated...flushing detector.
  • N3.75 FeII:
  • 18:22MST H2 and BrG give about 26k counts (sunset 18:40) but FeII too bright still.
  • N3.75 camera is way off! See 34, for example. Go to N1.8, click initialize, go back to N3.75. See 35 - better but still not centered. Back to N1.8 and then to N3.75. See 36. Looks like 35. Click initialize with N3.75 in place. 37 - looks like both 36 and 35. There are messages about "without limit switch"... i.e. Compensation mirror successfully iniitialized without limit switch.
  • flush detector
  • N3.75 and FeII flats 25k counts still at 18:31...
  • FeII just before sunset... 17k counts 41-50, flush_detector...(during which time I missed the 10k count level...) 7k-5k counts 52-71 some stars streaking through... increased DIT from 2 to 5sec. 8k counts: 72-81 and then back to 2sec for low illumination: 2k counts 82-91.
  • K: 5sec for 10k counts. 93-102. Try H at 2sec. 3sec H 104-113. K 5sec now gives 4k counts: 114-123. Try 2s J: 124. Only 3k counts. Back to H 3sec for low illum: 125-134 - about 2500 counts. 135 z at 3sec, 136 z at 10 sec 5k counts.
  • DONE with flats: Got some at FeII, K and H, albeit the high illumination did not have such high count levels, ~8k.
  • Need to set up a records page on the wiki to keep track

Pupil imaging, several rotations (two EL?)
  • Pointing at Zenith and parked --- flat field position
  • K filter, N1.8 PV to position IN (moves filter wheel 1 to PV lens), mirror, no mask. Rotator at 341, PA=0 Number 137 - there is something cutting across pupil opposite the swing arm. 138 is without K filter --- more light and less distinct image.
Note added later: This strip may have been from the crane. Clearance is small and if the telescope is parked at slightly beyond 90deg elevation, then it would see the crane. ?? How much does this affect LUCIFER flat fields. We would never have noticed this vignetting if not for putting in the pupil viewer, something which is not done on science observing nights.
  • 139 K filter and rotator at 251 deg. Still this obstruction is there, opposite swing arm.
  • 140 K filter and rotator at 161 deg. Again it is there.
  • K filter, N3.75 --- 141 (this looks like N1.8 camera and pupil viewer status is OUT now).
  • 142 K N1.8 PV IN (does N3.75 move pupil viewer out...?)
  • 143 K N1.8 PV IN ROTATOR at 161 deg. Looks like 140.
  • Pointing at 70 deg elevation
    • 144 K N1.8 PV IN ROT = 161. This vignetting opposite swing arm is gone now.
    • 145 ROT=251 deg.
    • 146 ROT=71 deg.
    • 147 ROT=341 deg.
  • Pointing at 40 deg elevation
    • 148 ROT=341 deg
    • 149 ROT=251 deg
    • 150 ROT=161 deg
    • 151 ROT=71 deg
  • I don't see any emission as bright as swing arm (cf Figure in LUCIFER manual) but I do see a dark crescent and changes in swing arm and secondary images as a function of rotator angle and elevation angle. We don't know what the bright band was seen at zenith - shutter doors are 100% open.

Correct initial collimation/pointing
  • 02:42 UT Slewing to ACT0333 for pointing correction. PaG+J 152. ptcorrbinoc: Current IE=-73, new = -83.5 and Current CA=63.6, new=61.5. 153 is image to verify motion. Looks like it moved. But is ptcorrbinoc the right task??
  • 154 is same star... run ptcorr on it. ptcorr does same thing - finds star in position and, as expected, does not move it much.
  • 02:57 UT Slewing to Persson star BS9122 with position/active. Good - seeing about 0.8".
  • While here, do K photometry using PhotomStdK.txt in /home/observer/LBTO/OlgaKuhn/ 2sec x 10 NDIT REL DETXY offsets. 70deg elevation. 156-160. But DIT/NDIT were reversed --- too bright. Repeating with 2sec and 10 NDIT (15k counts). 161-165.
  • 166-170 H and 171-175 J.

not done Open dome SW tests, if requested
  • Optical pointing/co-pointing/range balancing algorithm, also using IRTC

Run science-type dice5 pattern to check that things still work Several dither sequences done with RELATIVE offsets but none with ABSOLUTE
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/dice5_10x (~1 hr)
  • Any other source with a known ABS dither pattern is OK
  • Calculate rms/position, scale & rotation of pattern

Repeat check of dithering along slit (20x 2' dither) Not done
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/slitdither.acq
  • Align BS source to center of 2" longslit
  • Execute script: ~/LBTO/SciencePrep/slitdither.sci (~30 min)
  • Calculate AGw alignment to slit

Repeat measurement of focal plane scale, orientation DONE
  • Any low-latitude field with 20+ stars in USNO is OK
    • BS9122 field has many stars --- run /home/observer/LBTO/OlgaKuhn/imagepair using J instead of Ks since stars are bright. 8sec x 8 NDIT. 177 and 178.

Additional work on M1/M2 range-balancing collimation Not done

Field aberration data collection Not done

Check existing collimation model DONE: Looked OK for several Persson stars, but these were all around 50-60deg. When I slewed to a PhotStd near zenith (88 deg), saw it was out of focus, a donut.
  • Collimate at mid-EL
  • Check at low EL w/TRACK preset
  • Check at high EL w/TRACK preset
  • 03:52 Were collimated at BS9122 (last image was 178 J). N
  • Slewing to BS9143 (Position, Track) 179 J image at 53 deg elevation 8sec x 8 DIT. Image looks good.
  • Slewing to BS9139 (Position, Track) 180 J image at 62 deg elevation (not much change...) 8sec x 8 DIT.
  • FS125 is very high, elevation 88 deg. 4sec x 1 NDIT is 181 and 4sec x 16 NDIT is 182. Here collimation falls apart (focus).

* Photometry on FS125 at elev = 86 deg. J 184-188; H 189-193 about 17k counts; K F125 has K=194-198.

* Mask still does not move --- tried to insert POINTING mask, but immediately got an error message.

08:00 UT I've tried calling Walter and Peter for about 2.5 hours but no response yet and we switched to LBC - fast switch (~20 minutes), however Walter called just after that and we troubleshot the problem, which was in the Mask Config GUI (a check mark for "mask in use?" on NB 31 where there was no mask. Only a pair of LBC pupil images obtained - FPIA fit these (though one pupil was poorly fit).

09:34 UT Switching back to LUCIFER. Done at 09:55.

10:00 UT Pointing correction 200 and 201 and collimation on a Persson star nearby. We do not have very stable conditions... seeing is bouncing around 2.5" - 3" - it is terrible!

11:00 UT Practicing LUCIFER MOS acq.
  • Telluric acquisition: mask in turnout & star at center. Move star to 1013, 1042, then move mask in FPU and exit - which offsets and takes an image.
  • Make small adjustments viewing thru-slit image
  • Run telluric science script (once with not well-centered star, second time with centered star --- this second try may be useful to the test dithering between extreme slits of MOS mask (although the first two slits through which the star was observed are, as I recall, not too far apart).
  • MOS science. MOS out and acq runs two OFFSET images to allow sky-sub, then ends. Put mask in and take slit image. Use MOS ACQ GUI. Seeing so poor that only one reference star shows up. But cross-hatch boxes are over it, so WCS must be better than last month.
  • Tried RM7_spec_acq_telluric.HDF_K1.txt and another tell, plus science acq.
  • Field down at 35.
~12:00 Previous field was very low. Slewed to Persson star near zenith to checked IQ. Seeing is also jumping there, up to 3". Wind is near limits and we're in twilight. Closing.

Check existing pointing model Not done since after recovering from MOS GUI error, seeing was poor and not stable.
  • Use the "pointing" mask for focal plane measurements.

Photometric zero point and extinction monitoring Observed a few standards. Conditions were photometric.

Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
  • Check to make sure offsets still happen in a reasonable manner DONE --- OFFSETS are FAST, within 2sec, they look good

Improve map of AGw patrol field limits at high field angle Not Done
  • Requires set up work on a stone field

Test dithering between extreme slits on MOS mask + telluric Perhaps the test observation of a telluric will serve for this.
  • Manual needs 0.59836mm/" scale for the dXdY calculation
  • The how-to set up telluric observations needs updating for clarity.

Summary: We checked out LUCIFER and I practiced some MOS acq in poor seeing on one set of INAF scripts. Learned the importance of a tickmark. Think LUCI is ready for observing.

-- OlgaKuhn - 26 Mar 2011
Topic revision: r6 - 28 Mar 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback